OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. Show your work to receive full credit. y = x2 + 3x + 2 y = x2 + 2x + 1 Part 2: Using complete sentences, answer the following questions about the two quadratic equations above. Do the two quadratic equations have anything in common? If so, what? What makes y = x2 + 3x + 2 different from y = x2 + 2x + 1?

OpenStudy (anonymous):

Do you know how to factorize quadratic equations?

OpenStudy (anonymous):

no not really its an essay question for 10 points

OpenStudy (anonymous):

equations of the form \[Ax ^{2}+Bx+C =0\] can be factorized by the following steps I ) multiply A*C II ) then find the factors of the product such that the sum or difference of the the factors = B \[x^2+3x+2\] A= 1 B = 3 and C= 2 A*C = 2 factors of 2 are 1, 2 1+2 = 3 ( = B) III ) now split the middle term (3x) into x + 2x\[x^2+3x+2 = x^2 +x +2x+2 \]\[ x (x+1) +2 (x+1) = (x+1)(x+2)\] So the factors of the quadratic equation are (x+1) and (x+2)

OpenStudy (anonymous):

ok thank you so much @akitav

OpenStudy (anonymous):

Its been a pleasure :)

OpenStudy (anonymous):

work out the factors for the other equation and check whether they have anything in common.